OpenStudy (andriod09):

need help with a problem, in comments:

OpenStudy (andriod09):

help me find x please! \[2x^2-8=10\]

OpenStudy (anonymous):

X=3?

OpenStudy (andriod09):

how?

OpenStudy (anonymous):

x=3.01

OpenStudy (anonymous):

You take -8 to the rhs(right hand side) and then you divide 18 with 2 and then square root 9 to get 3.

OpenStudy (andriod09):

where'd you get the 9 from?

OpenStudy (andriod09):

@GoldPhenoix little help here please gold? Would mean alot

OpenStudy (goldphenoix):

So you have an equation: 2x^2 - 8 = 10. What is 2x^2 give you? Well another way of writing it is: 2x * 2x. That would give us 4x^2. So your equation should look like: 4x^2 - 8 = 10 To find x, you want to get rid of -8. To get rid of -8, you want to add 8 to both side -8 + 8 = 0 (Cancel out) 10 + 8 = 18 So your equation should look like: 4x^2 = 18 Now, what do we do next? There's many way. Hold on for a moment.

OpenStudy (andriod09):

\[x=9\] perhaps?

OpenStudy (goldphenoix):

Arg. I forgot how to do it. Sorry. =/

OpenStudy (goldphenoix):

But it's definitely not 9.

OpenStudy (andriod09):

ffs. What do you do to the \[4x^2\] to get rid of that exponent then?

OpenStudy (goldphenoix):

I remember now. So you have: 4x^2 = 18 You want to do reverse PEMDAS to find x, right? So you divide both side by 4. 4x^2 / 4 = x^2 18 / 4 = 4.5 So now your equation should look like: 4.5 = x^2

OpenStudy (goldphenoix):

To get rid of ^2, you want to square root both side.

OpenStudy (loser66):

\[2x^2 -8 =10\\+8~both~sides~ 2x^2 =18\\divide~both~side~by~2~you~got ~x^2 = 9\\squareroot~both~sides~x=3~and~x=-3\]

OpenStudy (anonymous):

^ is the way i was trying to explain

OpenStudy (anonymous):

no x=4 if use the qudnomial theory and apply is alongside with tan

OpenStudy (goldphenoix):

Ooh. My bad. T~T

OpenStudy (loser66):

ok